OpenStudy (anonymous):

Find the height of a cylinder with radius 25 cm and surface area 8953.54 cm2. A. 32.00 cm B. 16.00 cm C. 31.00 cm D. 62.00 cm

OpenStudy (callisto):

Surface area of the circle \(=2 \pi rh +2(\pi r^2)\) Put surface area=8953.54 and radius r = 25 into the formula \(8953.54 =2 \pi (25)h +2\pi( 25^2)\) \(8953.54- 2\pi( 25^2)=2 \pi (25)h\) \(8953.54- 2\pi( 25^2)=2 \pi (25)h\) \(5026.549=50 \pi h\) \(h=5026.549 / (50 \pi) = ...?\)